# Cointegration with Occasionally Binding Constraints

### v3

@@ 3.1 Nonlinearity and cointegration
Within the CKSVAR framework, a simple, non-trivial example of series that are $I^{\ast}(d)$ but not $I(d)$ is provided by the censored dynamic Tobit. When $\sum_{i=1}^{k}\phi_{i}^{+}=1$, this model has a unit root, and Bykhovskaya and Duffy (2024) show that $n^{-1/2}y_{\lfloor n\lambda\rfloor}^{+}\rightsquigarrow Y^{+}(\lambda)$ on $D[0,1]$, where $Y^{+}$ is a Brownian motion regulated (at zero) from below (their Theorem 3.2; see 3.3 below), and $\lVert\Delta y_{t}^{+}\rVert_{2+\delta_{u}}$ is uniformly bounded (their Lemma B.2). Thus $\Delta y_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{+}\sim I^{\ast}(1)$, even though, due to the nonlinearity in the model, neither series are $I(d)$ for any $d$.
@@ 3.2 The CKSVAR with unit roots
Our next step is to rewrite the CKSVAR, as in (LABEL:var-two-sided) or (LABEL:var-pm) above, in the form of a vector error-correction model (VECM). Define the autoregressive polynomials
@@
where $\Delta\coloneqq 1-L$ denotes the difference operator, and for clarity we note that $\Delta y_{t}^{+}=y_{t}^{+}-y_{t-1}^{+}$ (rather than being the positive part of $\Delta y_{t}$). In the case of a canonical CKSVAR, (LABEL:vecm-general) helpfully reduces to
@@
Just as in a linear (cointegrated) VAR, which corresponds to the special case of (LABEL:vecm) in which $\pi^{+}=\pi^{-}$ and $\gamma_{i}^{+}=\gamma_{i}^{-}$ for all $i\in\{1,\ldots,k-1\}$, the long-run dynamics will be governed by the matrix of coefficients on the lagged levels. More precisely, there are two such $p\times p$ matrices, $\Pi^{+}$ and $\Pi^{-}$, defined by
@@
on $D[0,1]$, where $U$ is Brownian motion in $\mathbb{R}^{p}$ with (positive definite) variance $\Sigma$.
@@
By the functional martingale central limit theorem, under the stated moment condition on $\{u_{t}\}$, a sufficient but not necessary condition for (LABEL:Unwkc) is that $\{u_{t}\}$ be a stationary and ergodic martingale difference sequence with $\Sigma\coloneqq\mathbb{E}u_{t}u_{t}^{\mathsf{T}}$. However, ERR also allows $\{u_{t}\}$ to be weakly dependent, such as if e.g. $u_{t}=\sum_{i=0}^{\infty}\Theta_{i}\eta_{t-i}$, where $\sum_{i=0}^{\infty}\lVert\Theta_{i}\rVert<\infty$ and $\{\eta_{t}\}$ is a mean zero, i.i.d. process with $\lVert\eta_{t}\rVert_{2+\delta_{u}}<\infty$.
In a linear VAR (i.e. where $\Phi^{+}(\lambda)=\Phi^{-}(\lambda)$) under CVAR the Granger–Johansen representation theorem (Johansen, 1995, Thm. 4.2 and Cor. 4.3) implies that $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}\sim I^{\ast}(1)$ and that there exists a full column rank matrix $\beta\in\mathbb{R}^{p\times r}$ such that $\beta^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$, where $r=r^{+}=r^{-}$. Thus $\{z_{t}\}$ is cointegrated in the sense of 3.2, with cointegrating space $\operatorname{sp}\beta$. Here DET specialises to $c\in\operatorname{sp}\Phi(1)$, preventing the model from generating any common deterministic trends between the series; it has the same effect more generally when $\Pi^{+}\neq\Pi^{-}$. (We shall relax this condition, so as to allow for deterministic trends, in 4.4 below.)
Case $r^{+}$ $r^{-}$ $\operatorname{rk}\Pi^{x}$ $\pi^{+}\in\operatorname{sp}\Pi^{x}$ $\pi^{-}\in\operatorname{sp}\Pi^{x}$
@@
While linear cointegration may occur in a CKSVAR, other phenomena are possible, depending on the ranks of $\Pi^{+}$, $\Pi^{-}$ and $\Pi^{x}$. Within the framework of $I^{\ast}(0)$ and $I^{\ast}(1)$ processes, as delimited by CVAR, there are three possibilities, each of which generate profoundly different trajectories for $\{y_{t}\}$. These are characterised in 3.1; in each case $r$ (without a superscript) is defined such that it is possible to regard the system as having a cointegrating rank of at least $r$ in each ‘regime’. Since, $\Pi^{+}$ and $\Pi^{-}$ differ by only their first column, $r^{+}$ and $r^{-}$ may differ by at most one, so that the case where $r^{+}\neq r^{-}$ may be identified with case (i) in the table without loss of generality. [^6] If we were to instead take $r^{+}=r^{-}+1$ in this case, the characteristic behaviour of the series in the positive and negative regimes, with $y_{t}^{+}\sim I^{\ast}(1)$ and $y_{t}^{-}\sim I^{\ast}(0)$, would now be reversed, but the key properties of the model would otherwise be unaltered.
@@ 3.3 Common trends and long-run trajectories
To develop some intuition for the properties of the model in these three cases, in advance of the representation theory developed in the next section, it is helpful to regard the long-run behaviour of the processes as being characterised by a space of common trends $\mathscr{M}$, defined as the set of non-stochastic (i.e. $u_{t}\coloneqq 0$, $\forall t$) steady state solutions to (LABEL:vecm-general):
@@
In a linear VAR, $\mathscr{M}=\ker\Pi$ is a $q$-dimensional linear subspace of $\mathbb{R}^{p}$, the orthogonal complement of which is the cointegrating space; and $Z$ is a $p$-dimensional Brownian motion with rank $q$ covariance matrix – i.e. it is a rank $q$ linear function of $U$ in (LABEL:Unwkc) above – taking values in $\mathscr{M}$. Whereas in the CKSVAR, $\mathscr{M}$ no longer need be a linear subspace, but is instead a linear cone [^7] Recall that a set $S\subset\mathbb{R}^{p}$ is termed a linear cone if $\lambda s\in S$ for every $s\in S$ and $\lambda\geq 0$, i.e. if $S$ is closed under multiplication by non-negative scalars. formed from the union of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, the vectors orthogonal to which are the cointegrating vectors $\beta^{+}$ and $\beta^{-}$ on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$ respectively (recall 3.2 above). While $Z$ remains a function of $U$, that function need not be linear: indeed, in addition to (linear) Brownian motions (BMs), any of the following nonlinear processes may also appear among the limiting stochastic trends generated by the CKSVAR. (For further discussion of regulated BMs, see Harrison, 1985, Ch. 1.)
@@
If $V$ is as in (i) or (ii), then $-V$ is respectively censored or regulated from above. Suppose now that $p\geq 1$, and let $G:\mathbb{R}\rightarrow\mathbb{R}^{p\times p}$ be a map that depends only on the sign of its argument, and is such that: (a) $h\coloneqq e_{1}^{\mathsf{T}}G(+1)=\mu e_{1}^{\mathsf{T}}G(-1)$ for some $\mu>0$; and (b) $w\mapsto G(h^{\mathsf{T}}w)w$ is continuous. Then the $p$-dimensional process $V$ is said to be a
@@
{adjustwidth}
-2cm-2cm Refer to caption Refer to caption Refer to caption Refer to caption
@@
Trajectories of these processes (denoted by $V$) in the univariate case ($p=1$), together with the realisation of the standard Brownian motion $W$ used to construct them, are plotted in 3.1. While both censored and regulated BMs are constrained to be positive, it is evident from panels (b) and (c) that there are important differences between them. For the former, the censoring does not feed back into the underlying dynamics, and $V$ spends long stretches at zero (while $W$ is negative); whereas for the latter, the $\sup_{\lambda^{\prime}\leq\lambda}[-W(\lambda^{\prime})]_{+}$ term continually reflects $V$ away from zero, so that $V$ spends relatively little time near zero. The kinked BM in panel (d) is constructed as
@@
For $p=2$, various possible shapes of $\mathscr{M}$ are illustrated graphically in Figures 3.23.4, along with matching example trajectories for $(y_{t},x_{t})$. In all figures, $(y_{t},x_{t})$ are generated by a canonical CKSVAR with $k=1$, $c=0$, $u_{t}\sim_{\textnormal{i.i.d.}}N[0,I_{2}]$, and $y_{0}=x_{0}=0$; the specification of the model is completed by specifying $\pi^{+}$, $\pi^{-}$ and $\Pi^{x}$ in (LABEL:vecm), the values of which are given in each panel. The associated cointegrating vectors (on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$) are the vectors orthogonal of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, while the form of the limiting processes $Z$ is suggested by the shape of the set $\mathscr{M}$, where it concentrates. The main qualitative features of the three cases, as enumerated in 3.1 above, are as follows.
@@ 3.3.1 Case (i): regulated cointegration
{adjustwidth}
-2cm-2cm Refer to caption Refer to caption
@@
The distinguishing characteristic of this case is that the common trends are restricted to the region where $y\geq 0$, so that $Y$ will always be a Brownian motion regulated (from below) at zero, even though $y_{t}$ itself may take negative values. In the case that $q=1$, as e.g. when $p=2$ and $r=1$ as depicted in 3.2, $X$ will also be a regulated process, i.e. we have cointegration where both processes share a common regulated stochastic trend. (When $q\geq 2$, $X$ will also depend on $q-1$ additional linear BMs.) Accordingly, a model configured as in case (i) would be most appropriate when $y_{t}$ appears to wander randomly above a threshold, and makes only brief sojourns below that threshold.
@@
with the associated CKSVAR for $(i_{t},\pi_{t})$ being as in (LABEL:cksvar-infldrift) above. In this first-order model, $\Pi^{\pm}=-\Phi^{\pm}(1)=\Phi_{1}^{\pm}-\Phi_{0}^{\pm}$, and thus it follows from (LABEL:cksvar-infldrift) that
@@
Suppose $\delta<0$, which in the context of 2.8 implies that past deviations from target indeed feed back into the central bank’s inflation target. Then unless $\mu=1$ (in which case the model is linear), $\varphi_{\mu}\neq\varphi_{1}$ and $\Pi^{-}$ has full rank. It follows that =$\mathscr{M}^{-}=\{0\}$, whereas since $\operatorname{rk}\Pi^{+}=1$, $\mathscr{M}^{+}$ is the ‘half’ subspace orthogonal to $\beta\coloneqq(1,-1)^{\mathsf{T}}$, as depicted in 3.2. It follows (via 4.1 below) that $i_{t}$ and $\pi_{t}$ are cointegrated, with cointegrating vector $\beta$, when $i_{t}$ is positive; but $I^{\ast}(0)$ when $i_{t}$ is negative; their common limiting stochastic trend is a regulated BM.
For the economics underlying this, note that (LABEL:is-pc21b) and (LABEL:taylor21b) imply that when the solution to the model has $i_{t}>0$, i.e. when the ZLB is not binding, monetary policy is able to fully achieve its objectives, in the sense that inflation is stabilised to within an i.i.d. error of its target, as
@@
As a consequence, (LABEL:infltarget) entails that $\overline{\pi}_{t}$ has a stochastic trend, which is inherited by both $\overline{\pi}_{t}$ and $i_{t}^{+}$ – in the latter case, because the equilibrium rate of interest implied by (LABEL:taylor21b) is equal to $\overline{\pi}_{t}$. Both $i_{t}^{+}$ and $\pi_{t}$ thus put the same loading on the common trend, whence $(1,-1)^{\mathsf{T}}$ is the cointegrating vector when $i_{t}>0$. On the other hand, when the solution to the model has $i_{t}<0$, i.e. when the ZLB binds, the lesser effectiveness of unconventional monetary policy ($\mu<1$) entails that policy is too contractionary, and so $\pi_{t}$ begins to drift below $\overline{\pi}_{t}$. However, via (LABEL:infltarget) this discrepancy raises the inflation target, and thereby raises the nominal rate of interest required to achieve target inflation. This feedback actually renders $\overline{\pi}_{t}$ as $I^{\ast}(0)$, and hence also $\pi_{t}$ and $i_{t}^{-}$. In a relatively short time, the solution to the model entails $i_{t}>0$ again, and thus the economy tends to spend relatively little time in the vicinity of the ZLB.

### v4

@@ 3.1 Nonlinearity and cointegration
Within the CKSVAR framework, a simple, non-trivial example of series that are $I^{\ast}(d)$ but not $I(d)$ is provided by the censored dynamic Tobit. When $\sum_{i=1}^{k}\phi_{i}^{+}=1$, this model has a unit root, and Bykhovskaya and Duffy (2024) show that $n^{-1/2}y_{\lfloor n\lambda\rfloor}^{+}\rightsquigarrow Y^{+}(\lambda)$ on $D[0,1]$, where $Y^{+}$ is a Brownian motion regulated (at zero) from below (their Theorem 3.2; see Definition 3.3 below), and $\lVert\Delta y_{t}^{+}\rVert_{2+\delta_{u}}$ is uniformly bounded (their Lemma B.2). Thus $\Delta y_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{+}\sim I^{\ast}(1)$, even though, due to the nonlinearity in the model, neither series are $I(d)$ for any $d$.
@@ 3.2 The CKSVAR with unit roots
Our next step is to rewrite the CKSVAR, as in (2.2) or (2.3) above, in the form of a vector error-correction model (VECM). Define the autoregressive polynomials
@@
where $\Delta\coloneqq 1-L$ denotes the difference operator, and for clarity we note that $\Delta y_{t}^{+}=y_{t}^{+}-y_{t-1}^{+}$ (rather than being the positive part of $\Delta y_{t}$). In the case of a canonical CKSVAR, (3.1) helpfully reduces to
@@
Just as in a linear (cointegrated) VAR, which corresponds to the special case of (3.2) in which $\pi^{+}=\pi^{-}$ and $\gamma_{i}^{+}=\gamma_{i}^{-}$ for all $i\in\{1,\ldots,k-1\}$, the long-run dynamics will be governed by the matrix of coefficients on the lagged levels. More precisely, there are two such $p\times p$ matrices, $\Pi^{+}$ and $\Pi^{-}$, defined by
@@
on $D[0,1]$, where $U$ is a Brownian motion in $\mathbb{R}^{p}$ with (positive definite) variance $\Sigma$.
@@
By the functional martingale central limit theorem, under the stated moment condition on $\{u_{t}\}$, a sufficient but not necessary condition for (3.4) is that $\{u_{t}\}$ be a stationary and ergodic martingale difference sequence with $\Sigma\coloneqq\mathbb{E}u_{t}u_{t}^{\mathsf{T}}$. However, ERR also allows $\{u_{t}\}$ to be weakly dependent, such as if e.g. $u_{t}=\sum_{i=0}^{\infty}\Theta_{i}\eta_{t-i}$, where $\sum_{i=0}^{\infty}\lVert\Theta_{i}\rVert<\infty$ and $\{\eta_{t}\}$ is a mean zero, i.i.d. process with $\lVert\eta_{t}\rVert_{2+\delta_{u}}<\infty$.
In a linear VAR (i.e. where $\Phi^{+}(\lambda)=\Phi^{-}(\lambda)$) under CVAR the Granger–Johansen representation theorem (Johansen, 1995, Thm. 4.2 and Cor. 4.3) implies that $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}\sim I^{\ast}(1)$ and that there exists a full column rank matrix $\beta\in\mathbb{R}^{p\times r}$ such that $\beta^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$, where $r=r^{+}=r^{-}$. Thus $\{z_{t}\}$ is cointegrated in the sense of Definition 3.2, with cointegrating space $\operatorname{sp}\beta$. Here DET specialises to $c\in\operatorname{sp}\Phi(1)$, preventing the model from generating any common deterministic trends between the series; it has the same effect more generally when $\Pi^{+}\neq\Pi^{-}$. (We shall relax this condition, so as to allow for deterministic trends, in Section 4.4 below.)
While linear cointegration may occur in a CKSVAR, other phenomena are possible, depending on the ranks of $\Pi^{+}$, $\Pi^{-}$ and $\Pi^{x}$. Within the framework of $I^{\ast}(0)$ and $I^{\ast}(1)$ processes, as delimited by CVAR, there are three possibilities, each of which generate profoundly different trajectories for $\{y_{t}\}$. These are characterised in Table 3.1; in each case $r$ (without a superscript) is defined such that it is possible to regard the system as having a cointegrating rank of at least $r$ in each ‘regime’. Since, $\Pi^{+}$ and $\Pi^{-}$ differ by only their first column, $r^{+}$ and $r^{-}$ may differ by at most one, so that the case where $r^{+}\neq r^{-}$ may be identified with case (i) in the table without loss of generality. [^6] If we were to instead take $r^{+}=r^{-}+1$ in this case, the characteristic behaviour of the series in the positive and negative regimes, with $y_{t}^{+}\sim I^{\ast}(1)$ and $y_{t}^{-}\sim I^{\ast}(0)$, would now be reversed, but the key properties of the model would otherwise be unaltered.
Case $^{+} rr^{-}$ $\operatorname{rk}\Pi^{x}$ $\pi^{+}\in\operatorname{sp}\Pi^{x}$ $\pi^{-}\in\operatorname{sp}\Pi^{x}$
@@ 3.3 Common trends and long-run trajectories
To develop some intuition for the properties of the model in these three cases, in advance of the representation theory developed in the next section, it is helpful to regard the long-run behaviour of the processes as being characterised by a space of common trends $\mathscr{M}$, defined as the set of non-stochastic (i.e. $u_{t}\coloneqq 0$, $\forall t$) steady state solutions to (3.1):
@@
In a linear VAR, $\mathscr{M}=\ker\Pi$ is a $q$-dimensional linear subspace of $\mathbb{R}^{p}$, the orthogonal complement of which is the cointegrating space; and $Z$ is a $p$-dimensional Brownian motion with rank $q$ covariance matrix – i.e. it is a rank $q$ linear function of $U$ in (3.4) above – taking values in $\mathscr{M}$. Whereas in the CKSVAR, $\mathscr{M}$ no longer need be a linear subspace, but is instead a linear cone [^7] Recall that a set $S\subset\mathbb{R}^{p}$ is termed a linear cone if $\lambda s\in S$ for every $s\in S$ and $\lambda\geq 0$, i.e. if $S$ is closed under multiplication by non-negative scalars. formed from the union of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, the vectors orthogonal to which are the cointegrating vectors $\beta^{+}$ and $\beta^{-}$ on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$ respectively (recall Definition 3.2 above). While $Z$ remains a function of $U$, that function need not be linear: indeed, in addition to (linear) Brownian motions (BMs), any of the following nonlinear processes may also appear among the limiting stochastic trends generated by the CKSVAR. (For further discussion of regulated BMs, see Harrison, 1985, Ch. 1.)
@@
If $V$ is as in (i) or (ii), then V-V is respectively censored or regulated from above. Suppose now that $p\geq 1$, and let $G:\mathbb{R}\rightarrow\mathbb{R}^{p\times p}$ be a map that depends only on the sign of its argument, and is such that: (a) $h\coloneqq e_{1}^{\mathsf{T}}G(+1)=\mu e_{1}^{\mathsf{T}}G(-1)$ for some $\mu>0$; and (b) $w\mapsto G(h^{\mathsf{T}}w)w$ is continuous. Then the $p$-dimensional process $V$ is said to be a
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Trajectories of these processes (denoted by $V$) in the univariate case ($p=1$), together with the realisation of the standard Brownian motion $W$ used to construct them, are plotted in Figure 3.1. While both censored and regulated BMs are constrained to be positive, it is evident from panels (b) and (c) that there are important differences between them. For the former, the censoring does not feed back into the underlying dynamics, and $V$ spends long stretches at zero (while $W$ is negative); whereas for the latter, the $\sup_{\lambda^{\prime}\leq\lambda}[-W(\lambda^{\prime})]_{+}$ term continually reflects $V$ away from zero, so that $V$ spends relatively little time near zero. The kinked BM in panel (d) is constructed as
@@
For $p=2$, various possible shapes of $\mathscr{M}$ are illustrated graphically in Figures 3.23.4, along with matching example trajectories for $(y_{t},x_{t})$. In all figures, $(y_{t},x_{t})$ are generated by a canonical CKSVAR with $k=1$, $c=0$, $u_{t}\sim_{\textnormal{i.i.d.}}N[0,I_{2}]$, and $y_{0}=x_{0}=0$; the specification of the model is completed by specifying $\pi^{+}$, $\pi^{-}$ and $\Pi^{x}$ in (3.2), the values of which are given in each panel. The associated cointegrating vectors (on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$) are the vectors orthogonal of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, while the form of the limiting processes $Z$ is suggested by the shape of the set $\mathscr{M}$, where it concentrates. The main qualitative features of the three cases, as enumerated in Table 3.1 above, are as follows.
@@ 3.3.1 Case (i): regulated cointegration
Refer to caption Refer to caption
@@
The distinguishing characteristic of this case is that the common trends are restricted to the region where $y\geq 0$, so that $Y$ will always be a Brownian motion regulated (from below) at zero, even though $y_{t}$ itself may take negative values. In the case that $q=1$, as e.g. when $p=2$ and $r=1$ as depicted in Figure 3.2, $X$ will also be a regulated process, i.e. we have cointegration where both processes share a common regulated stochastic trend. (When $q\geq 2$, $X$ will also depend on q1q-1 additional linear BMs.) Accordingly, a model configured as in case (i) would be most appropriate when $y_{t}$ appears to wander randomly above a threshold, and makes only brief sojourns below that threshold.
@@
with the associated CKSVAR for $(i_{t},\pi_{t})$ being as in (2.9) above. In this first-order model, $\Pi^{\pm}=-\Phi^{\pm}(1)=\Phi_{1}^{\pm}-\Phi_{0}^{\pm}$, and thus it follows from (2.9) that
@@
Suppose $\delta<0$, which in the context of 2.8 implies that past deviations from target indeed feed back into the central bank’s inflation target. Then unless $\mu=1$ (in which case the model is linear), $\varphi_{\mu}\neq\varphi_{1}$ and $\Pi^{-}$ has full rank. It follows that $\mathscr{M}^{-}=\{0\}$, whereas since $\operatorname{rk}\Pi^{+}=1$, $\mathscr{M}^{+}$ is the ‘half’ subspace orthogonal to $\beta\coloneqq(1,-1)^{\mathsf{T}}$, as depicted in Figure 3.2. It follows (via Theorem 4.1 below) that $i_{t}$ and $\pi_{t}$ are cointegrated, with cointegrating vector $\beta$, when $i_{t}$ is positive; but $I^{\ast}(0)$ when $i_{t}$ is negative; their common limiting stochastic trend is a regulated BM.
For the economics underlying this, note that (2.4b) and (2.5b) imply that when the solution to the model has $i_{t}>0$, i.e. when the ZLB is not binding, monetary policy is able to fully achieve its objectives, in the sense that inflation is stabilised to within an i.i.d. error of its target, as
@@
As a consequence, (2.8) entails that $\overline{\pi}_{t}$ has a stochastic trend, which is inherited by both $\overline{\pi}_{t}$ and $i_{t}^{+}$ – in the latter case, because the equilibrium rate of interest implied by (2.5b) is equal to $\overline{\pi}_{t}$. Both $i_{t}^{+}$ and $\pi_{t}$ thus put the same loading on the common trend, whence $(1,-1)^{\mathsf{T}}$ is the cointegrating vector when $i_{t}>0$. On the other hand, when the solution to the model has $i_{t}<0$, i.e. when the ZLB binds, the lesser effectiveness of unconventional monetary policy ($\mu<1$) entails that policy is too contractionary, and so $\pi_{t}$ begins to drift below $\overline{\pi}_{t}$. However, via (2.8) this discrepancy raises the inflation target, and thereby raises the nominal rate of interest required to achieve target inflation. This feedback actually renders $\overline{\pi}_{t}$ as $I^{\ast}(0)$, and hence also $\pi_{t}$ and $i_{t}^{-}$. In a relatively short time, the solution to the model entails $i_{t}>0$ again, and thus the economy tends to spend relatively little time in the vicinity of the ZLB.